OpenStudy (anonymous):

For the fibonacci numbers fn prove that the gcd(fn, fn+1)=1 for all n greater than or equal to 1. Do I need to maybe do induction?

OpenStudy (anonymous):

lol fibonacci = induction

OpenStudy (anonymous):

\[gcd(f_{n+1}, f_{n+1})=gcd(f_{n+1},f_{n+1}+f_{n+2})\] is probably a good place to start, since \(f_n\) is the fibonacci sequence

OpenStudy (anonymous):

damn typo!

OpenStudy (anonymous):

\[gcd(f_{n+1}, f_{n+2})=gcd(f_{n+1},f_{n+1}+f_{n+2})\] is more like it

OpenStudy (anonymous):

damn another tyopo i must be tired let me try again

OpenStudy (anonymous):

\[gcd(f_{n+1}, f_{n+2})=gcd(f_{n+1},f_{n}+f_{n+1})\] whew that is what i was trying to point out that is because \(f_{n+2}=f_n+f_{n+1}\)

OpenStudy (anonymous):

it's ok at least you are trying to help hahaha! Thanks for all of your help tonight